Write the equation of a line that passes through the points (7,-1) and (3,-3) in slope-intercept form

How do i do this???
1
Expert's answer
2012-01-05T10:50:50-0500
y=kx+b so we know that -1=7k+b -3=3k+b so k=0.5, b=-4.5 y=0.5x-4.5